# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v4

@@ 5 Conclusion
In this paper, we have explored the effective ways to prompt SAM with texts and demonstrate the importance of using the Multimodal Encoder with early fusion and multimodal inputs, i.e., text prompts and input images. To this end, we propose EVF-SAM, which establishes a new and simple path for extending SAMs’ text-prompted segmentation abilities with the off-the-shelf foundation models. We conduct experiments on the referring expression segmentation (RES) tasks with various benchmarks to evaluate the performance of text-prompted SAM. Experimental results showcase that our EVF-SAM achieves state-of-the-art performance for segmenting objects with referring texts on RefCOCO/+/g benchmarks, outperforming recent approaches based on Large Language Models with huge numbers of parameters. Moreover, experiments prove that (1) a multimodal encoder with input text and image and (2) the early fusion between image and text do matter more for prompting SAM than vanilla text encoders or Large Language Models. We hope this study and experiments can bring new ideas or insights to inspire future research on prompting SAM with texts.
Acknowledgments
This work was partially supported by the National Science and Technology Major Project under Grant No. 2023YFF0905400 and the National Natural Science Foundation of China (NSFC) under Grant No. 62276108. We thank Bo Jiang for his helpful feedback on the draft.
Appendix A Appendix
A.1 Qualitative Results
In this section, we mainly visualize the qualitative results on RefCOCO val and RefCOCOg val datasets, as shown in Fig. 6 and Fig. 7, respectively. Moreover, we compare the qualitative results of different ways to prompt SAM with texts: (1) our proposed EVF-SAM, (2) SAM with LLM (LISA (Lai et al., 2023)), and (3) SAM with a CLIP text encoder implemented in this paper (suggested by (Kirillov et al., 2023), which are based on the same SAM-Huge model. The qualitative results can demonstrate the superiority of the proposed EVF-SAM.
Visualizations on RefCOCOg. Fig. 7 illustrates the qualitative comparisons on the RefCOCOg val, which aims to segment objects with long expression texts. The SAM with a vanilla CLIP text encoder produces inferior segmentation results given the long-expression texts. However, the proposed EVF-SAM outperforms LISA when using long expressions, even though LISA adopts LLaMA-7B (Touvron et al., 2023) to understand the instructions and generate prompt embeddings, showcasing that the lightweight vision-language models can understand complex expressions. In addition, the proposed EVF-SAM can also understand the texts or expressions towards spatial locations, such as ‘the umbrella closest to the camera’.
A.2 Training EVF-SAM with Multi-Tasks
To further enhance the generic capability of our EVF-SAM, we propose to implement multi-task training. Based on the experiments that show the performance degradation when simply including extra segmentation data, we explore ways to make our EVF-SAM gain from extra data.
Mixed training with semantic segmentation. We introduce some extra semantic segmentation datasets (ADE20K (Zhou et al., 2017b), Mapillary (Neuhold et al., 2017)) to proceed with joint training. We do not include COCO-Stuff (Caesar et al., 2018) to avoid data leakage with RefCOCO/+/g. It can be seen in Tab. 7 that the performance on RefCOCO+ and ADE20K gains, indicating the effectiveness of including extra data to enhance the generic capability. However, the evaluation metrics of RefCOCO and RefCOCOg decrease when simply including extra semantic segmentation data. We owe this phenomenon to semantic conflict (Liu et al., 2023e).
@@ A.2 Training EVF-SAM with Multi-Tasks
Table 7: Results of adding extra semantic data. means zero-shot results. The reported ADE20K results are evaluated on the validation set using the cIoU metric.
@@
Unified training with multi-task datasets. To solve the semantic conflict mentioned above, we propose several pre-process strategies for datasets of different distributions. We will open-source related codes in our project page.
- Instance-level data: We apply Objects365 (Shao et al., 2019) to extend RES data. Specifically, (a) for each image, we exclude categories with more than one instance to avoid ambiguity problem. (b) we employ SAM-2 (Ravi et al., 2024) to automatically annotate masks according to the selected ground-truth bounding boxes. The remaining annotations maintain a rich amount thanks to the dense annotation of Objects365 (Shao et al., 2019). We obtain 524K images (of original 600K images) with 1.8M annotations (of original 10M annotations). The mask quality from automatic annotation is fine thanks to the accurate ground-truth from Objects365 (Shao et al., 2019) and the powerful segmentation capability of SAM-2 Ravi et al. (2024). Besides, the remaining annotations are valuable for addressing long-tail problems because those excluded annotations often belong to head categories.
- Semantic-level data: We introduce ADE20K (Zhou et al., 2017a; 2019) to broaden multi-task capability. We construct a special token ‘[semantic]’ and input ‘[semantic] {category}’. The special token would not be limited to common grammar so it is helpful to avoid semantic conflict.
- Part-level data: To enable the model to segment parts of objects, we introduce PartImageNet (He et al., 2022), HumanParsing (Liang et al., 2015a; b) and PASCAL-Part (Chen et al., 2014) to train our model. For semantic-level annotated datasets, i.e., HumanParsing, we implement the same strategy as ADE20K. Exceptionally, we align the definition of ‘left’ and ‘right’ with RES datasets (e.g., RefCOCO). For instance-level annotated datasets, i.e., PartImageNet and PASCAL-Part, we merge instance masks of the same category to convert the dataset to semantic-level. Then, the same strategy as ADE20K is implemented.
By combining those datasets, we observe a significant performance gain of 1.0 cIoU on the average metric, as shown in Tab. 2. Moreover, our model is able to proceed with multiple tasks like part-segmentation and semantic-level segmentation.

### v5

@@ 5 Conclusion
In this paper, we explore the effective ways to prompt SAM with texts and demonstrate the importance of using the Multimodal Encoder with early-fusion and multimodal inputs. To this end, we propose EVF-SAM, which establishes a new and simple path for extending SAMs’ text-prompted segmentation abilities with the off-the-shelf foundation models. We conduct experiments on the Referring Expression Segmentation (RES) tasks with various benchmarks to evaluate the performance of text-prompted SAM. Experimental results showcase that our EVF-SAM achieves state-of-the-art performance for segmenting objects with referring texts on RefCOCO/+/g benchmarks, outperforming recent approaches based on Large Language Models with huge numbers of parameters. Moreover, we propose carefully-designed data strategies for balancing conflicts from data of different distributions. We construct a harmonious hybrid dataset, leading to an uni-model that performs well on various segmentation tasks. We hope this study can bring new ideas or insights to inspire future research on prompting SAM with texts.
Supplementary Material
Appendix A Language-guided Multi-task Segmentation.
Due to the limitation of Referring Expression Segmentation (RES) data in aspects of both quantity (80k) and quality (mask accuracy, caption diversity), employing data from other sub-tasks of segmentation (e.g., semantic segmentation) become popular among recent works (e.g., LISA [25], u-LLaVA [65]). If properly implemented, additional data can not only improve RES performance, but also extend the segmentation granularities of RES (i.e., semantic-level segmentation, part-level segmentation, introduced in Sec. A.3 and Sec. A.4). However, existing works usually apply a rough manner, leading to weak or even negative influence on models’ RES performance. Moreover, they are weak at other granularites of RES.
In this section, we firstly analyze the challenges of employing data from other sub-tasks of segmentation (e.g., semantic segmentation), then we present our detailed recipe of our data strategies to building the hybrid dataset. Finally we introduce two tasks called ”Semantic-level Referring Expression Segmentation” and ”Part-level Referring Expression Segmentation” to evaluate the models’ performance at different granularites of RES.
Table 7: Explanation to semantic conflict. The top row contains the source image and annotated mask in the ADE20K [77, 78] dataset. The middle row presents the comparison of templates used by different works to apply semantic data into Referring Expression Segmentation, where red fonts represent the wrong description, while the green fonts represent the right description. The bottom row contains ideal caption. Previous works either prompt with the raw category name without modifications or apply a fixed template, leading to unaligned descriptions in some cases. We use special token ‘[semantic]’ as the task identifier, overcoming the misalignment problem.
Image & mask [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Raw [25] cat cat man
Fixed template [39] all the cats all the cats all the mans
Special token (ours) [semantic] cat [semantic] cat [semantic] man
Ideal caption all the cats/cats/… the cat/cat/… all the men/…
A.1 Challenges.
Existing segmentation data can be generally separated to the semantic-level data and the instance-level data. When applying these data for Referring Expression Segmentation (RES), Semantic conflict will arise in the semantic-data and ambiguity problem will arise in the instance-level data.
Semantic conflict: The semantic conflict, as discussed in UniLSeg [39], mainly exhibits between semantic segmentation datasets and RES datasets. Tab. 7 presents several representative cases. The semantic segmentation datasets only provide the category name as annotation, letting go of the information regarding to instance number. However, in RES tasks, the description should explicitly align with the related region. As can be seen in the first case in Tab. 7, ‘cat’ does not align with all three cats in the image. Some works (i.e., [39]) propose using fixed template like ‘all the category’ to avoid the semantic conflict, while introducing additional misalignment. The over-description problem can be seen in the second case in Tab. 7. ‘All the cats’ is used to describe the single cat in the image. Furthermore, difficulty of handling plural forms can be seen in the third case in Tab. 7. The fixed template would fail at categories like ‘bread’, ‘man’. The experiment in Sec. B.2 shows how semantic conflict affects performance.
Ambiguity problem: The ambiguity problem mainly exhibits between instance-level dataset and RES datasets. In instance-level datasets, multiple objects may align to the same category. However, in RES datasets, the correspondence is one-to-one. Directly using instance segmentation data to train an RES model will result in inconsistent target assignments, leading to unstable training.
A.2 Unified training with multi-task datasets.
To solve the problems mentioned above, we propose data strategies for various datasets. Specifically, we propose filter/merging principles for instance-level data and introduce special token ‘[semantic]’ to semantic-level data. We will open-source related codes on our project page.
Instance-level data: We include Objects365 [51] to extend RES data. Specifically, (a) for each image, we exclude categories with more than one instance to avoid ambiguity problems. (b) we employ SAM-2 [48] to automatically annotate masks according to the selected ground-truth bounding boxes. Despite filtering, a substantial number of annotations remain, owing to the high density of annotations in Objects365 [51]. From the original 600K images and 10M annotations, we retain 524K images and 1.8M annotations. The mask quality from automatic annotation is high due to the accurate ground-truth from Objects365 [51] and the powerful segmentation capability of SAM-2 [48]. Notably, the remaining annotations are valuable for addressing long-tail problems because those excluded annotations predominantly represent head categories.
Semantic-level data: We introduce ADE20K [77, 78] to broaden RES models’ capability of various granularities. We introduce a special token ‘[semantic]’ and input ‘[semantic] {category}’ to RES models. The special token would not be limited to common grammar so it is helpful to avoid semantic conflict. As shown in Tab. 7, the reconstructed descriptions with our special token have no conflict with the mask annotation.
Part-level data: To enable the model to segment parts of objects, we introduce PartImageNet [12], HumanParsing [31, 32] and PASCAL-Part [5] to train our model. For datasets annotated in semantic-level, i.e., HumanParsing, we implement the same strategy as ADE20K. Exceptionally, we align the definition of ‘left’ and ‘right’ with RES datasets (e.g., RefCOCO). For datasets annotated in instance-level, i.e., PartImageNet and PASCAL-Part, we merge instance masks of the same category to convert the dataset to semantic-level. Then, the same strategy as ADE20K is implemented.
By combining those datasets, we observe a significant performance gain of 1.0 cIoU in average on RefCOCO/+/g [21, 72, 42, 41], as shown in Tab. 1. Moreover, our model is empowered the capability of RES of various granularities (i.e., semantic-level segmentation, part-level segmentation, introduced in Sec. A.3 and Sec. A.4).
A.3 Semantic-level RES
In this section, we will make a detailed explanation of our proposed new task ‘Semantic-level Referring Expression Segmentation’, present the metric comparison, and provide some visualizations.
Definition. Basic RES tasks only focus on the instance-level segmentation. A special prompt only refers to a single object in the image. Trained only on RefCOCO/+/g [21, 72, 42, 41] datasets, the model cannot correctly respond to more complex prompts (e.g., multiple objects, stuff categories).
To this end, we propose ‘Semantic-level RES’, aiming at improving RES models’ capability of segmenting semantic-level objects. We make perfect use of existing semantic segmentation datasets. The annotated category is managed as text prompt of the RES model, while the semantic mask corresponding to the category is managed as prediction target. Notably, we include stuff categories during training and evaluation.
‘Semantic-level RES’ differs from semantic segmentation in that we manage object category as model input, not the output. ‘Semantic-level RES’ differs from open vocabulary segmentation in that this is an in-domain task.
Metrics on Semantic-level RES. Being a sub-task of Referring Segmentation, we take cIoU and gIoU as metrics following the RES setting. This metric examines the competence of a model to segment multi-objects and stuff categories.
We compare our model with LISA [25] and u-LLaVA [65], both of which also include ADE20K [77, 78] data during training. As shown in Tab. 8, our model demonstrates a significant advantage over previous methods in semantic-level RES. This superior performance is attributed to our novel approach to resolving the semantic conflict (i.e., using the special token ‘[semantic]’). The solution of semantic conflict greatly improves model’s capability of segmenting background classes (e.g., ‘sky’, ‘sea’, or simply ‘background’) and segmenting multi-objects (e.g., ‘people’, ‘flowers’).
Visualizations on ADE20K. Fig. 6 shows the visualization of ‘Semantic-level Referring Segmentation’ task on ADE20K val. Given an semantic annotated category as text prompt for RES models, the models are to predict semantic masks containing all pixels corresponding to the category. Visualization results are provided by our EVF-SAM. Trained on Semantic-level Referring Segmentation tasks, EVF-SAM is able to segment multiple objects (e.g., drawings, chairs) or background categories (e.g., floor, ground).
Figure 6: Visualization Results on ADE20K Val. We add a special token ‘[semantic]’ in front of category name to construct the referring prompt, as written below each image. We use red boxes to highlight the targeted object and use blue masks to visualize the segmentation result. Upper three images showcase EVF-SAM’s capability at multi-object RES. Below three images showcase EVF-SAM’s capability at stuff-object RES.
Table 8: Semantic-level Referring Segmentation results. * denotes zero-shot results. EVF-SAM outperforms previous methods that include ADE20K [77, 78] during training, indicating the effectiveness of our proposed special token ‘[semantic]’.
Method ADE20K [77, 78] Pascal VOC [11]
gIou cIoU gIou cIoU
LISA-7B 49.0 67.4 53.6* 46.7*
LISA-7B-explanatory 55.4 70.8 56.2* 47.0*
u-LLaVA 42.1 75.3 57.5 44.8
EVF-SAM 64.5 76.8 78.2* 81.8*
A.4 Part-level RES
In this section, we will make a detailed explanation of our proposed new task ‘Part-level RES’, present the metric comparison, and provide some visualizations.
Definition Similar to Semantic-level RES, we further propose ‘Part-level RES’, aiming at improving RES models’ capability of segmenting part of objects. We make perfect use of existing part segmentation datasets. The annotated category is managed as text prompt of the RES model, while the semantic mask corresponding to the category is managed as prediction target.
Metrics on Part-level RES We evaluate part-level RES on Pascal-Part dataset. It is worth noting that we merge the instance-level annotations to semantic-level annotations to avoid ambiguity problems, as illustrated in Sec. 3.3. We still following the RES settings to evaluate cIoU and gIoU as validation metrics. The metrics are shown in Tab. 9. We compare our EVF-SAM with LISA [25], which also includes Pascal-Part [5] dataset during training. The experiment results show that our EVF-SAM outperforms LISA at part-level RES task by a huge margin.
Visualizations on Pascal-Part. Fig. 7 shows the visualization of ‘Part-level Referring Expression Segmentation’ on Pascal-Part val.
Figure 7: Visualization Results on Pascal-Part Val. We add a special token ‘[semantic]’ in front of category name to construct the referring prompt, as written below each image. We use red boxes to highlight the targeted object and use blue masks to visualize the segmentation result. Upper three images showcase EVF-SAM’s capability at multi-part RES. Below three images showcase EVF-SAM’s capability at broad-part RES.
Table 9: Part-level Referring Expression Segmentation results. We present evaluation metrics on Pascal-Part and compare the results with LISA, which also includes Pascal-Part during training. Our EVF-SAM outperforms LISA by a large margin.
Method gIoU cIoU
LISA 21.2 27.0
EVF-SAM 53.7 71.4
Appendix B Additional Experiments
B.1 Reasoning segmentation.
ReasonSeg [25] is a Referring Expression Segmentation dataset targeting long and indirect descriptions. It comprises 239 training images, 200 validation images and 779 test images. Text annotations of ReasonSeg are posed in an instructional format, such as ‘What is the food with the most Vitamin C in this image?’. Considering the limited amount of data and the risk of over-fitting, we process evaluation using both the zero-shot setting and the finetuning setting. The zero-shot performance on ReasonSeg of our EVF-SAM surpasses LLaVA-free methods (e.g., OVSeg, SEEM), but falls behind methods with LLaVA1.1 [36]. When we finetune our EVF-SAM with 239 ReasonSeg training samples, we observe significant improvement in reasoning performance, outperforming LISA-7B-LLaVA1.1 [25] However, finetuned EVF-SAM still falls behind LISA with LLaVA1.5 [37].
We analyze LISA’s superior performance on ReasonSeg compared to EVF-SAM, attributing it to several key factors:
  • LISA’s reasoning abilities significantly improved when transitioning from LLaVA 1.1 to LLaVA 1.5. This enhancement stems from LLaVA 1.5’s incorporation of reasoning-focused Visual Question Answering (VQA) datasets (e.g., CLEVR [19], GQA [16]), which effectively transfer reasoning knowledge to ReasonSeg. However, EVF-SAM’s foundation model (i.e., BEIT-3 [59]) is only pre-trained using Masked Language Modeling (MLM) recipe.
  • LLM-free methods (e.g., EVF-SAM) are absent of instruction tuning, hindering their ability to accurately interpret the QA formatted data.
Table 10: Comparision of segmentation capability on ReasonSeg [25] dataset. Our EVF-SAM outperforms previous methods on both cIoU and gIoU metrics.
Method Finetune? val test
gIoU cIoU gIoU cIoU
OVSeg 28.5 18.6 26.1 20.8
X-Decoder 22.6 17.9 21.7 16.3
SEEM 25.5 21.2 24.3 18.7
LISA-7B-LLaVA1.1 44.4 46.0 36.8 34.1
LISA-7B-LLaVA1.1 52.9 54.0 36.8 34.1
LISA-7B-LLaVA1.5 53.6 52.3 48.7 48.8
LISA-7B-LLaVA1.5 61.3 62.9 55.6 56.9
LISA-13B-LLaVA1.5 65.0 72.9 61.3 62.2
EVF-SAM 38.3 30.0 38.9 32.6
EVF-SAM 58.7 55.7 55.1 50.5
B.2 Experiments of simply adding extra data.
We tried to introduce some extra semantic segmentation datasets (ADE20K [77, 78], Mapillary [43]) to proceed with joint training. We do not include COCO-Stuff [3] to avoid data leakage with RefCOCO/+/g. We report comparison results in Tab. 11. From the experimental results: (a) We find that several metrics gain when introducing extra semantic segmentation data (e.g., RefCOCO+/testB). This may results from the analogous data composition between RefCOCO+ and ADE20K, indicating the feasibility of introducing data of other segmentation sub-task to boost RES model. (b) We find that the Semantic-level Referring Segmentation capability improves a lot when including extra semantic segmentation data, indicating the necessity of building multi-task training. (c) However, We observe performance degradation on various metrics (e.g., RefCOCO/val, RefCOCO/testA). This is due to the difference distribution between datasets from RES and other segmentation sub-tasks.
@@ Appendix B Additional Experiments
Table 11: Results of adding extra semantic data. means zero-shot results. The reported ADE20K results are evaluated on Semantic-level Referring Segmentation tasks proposed in main text.
@@
B.3 Multimodal feature representation.
[…]
- 83.5
- 83.7
- 83.5
Concat 83.2
In Tab. 12, we explore the effects of using different multimodal features representations as prompts for SAM. Specifically, we adopt different outputs of the Multimodal Encoder: (a) the image [CLS] token, (b) the AvgPool over image tokens, and (c) the text [CLS] token. Tab. 12 shows that using image [CLS] token is more effective while combining image and text tokens through concatenation leads to a performance drop.
Appendix C Qualitative Results
In this section, we mainly visualize the qualitative results on RefCOCO val and RefCOCOg val datasets, as shown in Fig. 9 and Fig. 10, respectively. Moreover, we compare the qualitative results of different ways to prompt SAM with texts: (1) our proposed EVF-SAM, (2) SAM with LLM (LISA [25]), and (3) SAM with a CLIP text encoder implemented in this paper (suggested by [24], which are based on the same SAM-Huge model. The qualitative results can demonstrate the superiority of the proposed EVF-SAM.
Visualizations on RefCOCOg. Fig. 10 illustrates the qualitative comparisons on the RefCOCOg val, which aims to segment objects with long expression texts. The SAM with a vanilla CLIP text encoder produces inferior segmentation results given the long-expression texts. However, the proposed EVF-SAM outperforms LISA when using long expressions, even though LISA adopts LLaMA-7B [57] to understand the instructions and generate prompt embeddings, showcasing that the lightweight vision-language models can understand complex expressions. In addition, the proposed EVF-SAM can also understand the texts or expressions towards spatial locations, such as ‘the umbrella closest to the camera’.